Solve the following equations:
p/4 = 5.................

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

We have to multiply both sides of the equation by 4,
Then, we get,
= p/4 × 4 = 5 × 4
= p = 20

Thanks